OpenStudy (anonymous):

Data Set 1 has a mean of 120 and a MAD of 4. Data Set 2 has a mean of 132 and a MAD of 3. What can you conclude about the two distributions? Choose exactly two answers that are correct. A. The means-to-MAD ratio is 3. B. The means-to-MAD ratio is 4. C. The distributions are somewhat similar. D. The distributions are different.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Here_to_Help15 @HelpBlahBlahBlah

OpenStudy (anonymous):

can anyone help?

OpenStudy (michele_laino):

I think that your distributions are different each other, since the discrepancy, namely the difference 132-120=12 is greater than the MAD of both distributions

OpenStudy (anonymous):

so D and what other answer?

OpenStudy (michele_laino):

yes! nevertheless ask to another helper also

OpenStudy (anonymous):

@Michele_Laino there has to be 2 answers

OpenStudy (michele_laino):

here we have: mean to MAD ratio is 120/4 = 30 for first distribution

OpenStudy (anonymous):

so d and b?

OpenStudy (michele_laino):

b. no, since the mean-to-MAD ratio is 132/3 = 44

OpenStudy (anonymous):

b and C?

OpenStudy (anonymous):

*D and C

OpenStudy (michele_laino):

please wait I try to ask to another helper

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

@dan815 please can you help here?

OpenStudy (michele_laino):

@e.mccormick can you help here please?

OpenStudy (michele_laino):

@triciaal can you hel here please?

OpenStudy (triciaal):

sorry no by the way mean to MAD is 30 not 3

OpenStudy (anonymous):

so D and A

OpenStudy (anonymous):

It's D and A